### v4

# Root Cause Analysis of Outliers with Missing Structural Knowledge

Abstract
The goal of Root Cause Analysis (RCA) is to explain why an anomaly occurred by identifying where the fault originated. Several recent works model the anomalous event as resulting from a change in the causal mechanism at the root cause, i.e., as a soft intervention. RCA is then the task of identifying which causal mechanism changed. In real-world applications, one often has either few or only a single sample from the post-intervention distribution: a severe limitation for most methods, which assume one knows or can estimate the distribution. However, even those that do not are statistically ill-posed due to the need to probe regression models in regions of low probability density. In this paper, we propose simple, efficient methods to overcome both difficulties in the case where there is a single root cause and the causal graph is a polytree. When one knows the causal graph, we give guarantees for a traversal algorithm that requires only marginal anomaly scores and does not depend on specifying an arbitrary anomaly score cut-off. When one does not know the causal graph, we show that the heuristic of identifying root causes as the variables with the highest marginal anomaly scores is causally justified. To this end, we prove that anomalies with small scores are unlikely to cause those with larger scores in polytrees and give upper bounds for the likelihood of causal pathways with non-monotonic anomaly scores.
1 Introduction
When an anomaly occurs, it is essential to identify its root cause so that action can be taken to resolve it. As such, Root Cause Analysis (RCA), has become a rich and rapidly developing area of research, with wide ranging applications from meteorology Wibisono et al. (2021), monitoring health Strobl and Lasko (2023); Strobl (2024), industrial fabrication Susto et al. (2017), fraud detection Vanhoeyveld et al. (2020), credit scoring Das et al. (2023), cloud applications Wang et al. (2023); Lin et al. (2024); Liu et al. (2021); Ikram et al. (2022); Shan et al. (2019); Ma et al. (2020) and more Hawkins (1980); Chandola et al. (2009); Aggarwal (2017); Blázquez-García et al. (2021); Akoglu (2021). Although many previous works have adopted a purely statistical approach to RCA Knorr and Ng (1999); Micenková et al. (2013); Liu et al. (2017); Macha and Akoglu (2018); Gupta et al. (2019), relying on correlations, the fact that the goal is to take action distinguishes it as a necessarily causal problem. The predominant choice in causal literature is to model the observed anomaly as being the result of a change in the causal mechanism at the root cause, i.e., as a soft intervention. RCA can then be broadly understood as the task of identifying the variable whose mechanism changed. Budhathoki et al. Budhathoki et al. (2022) give a formalisation of RCA as a quantitative contribution analysis based on counterfactuals – asking, “would the observed anomaly still be anomalous had this mechanism been as normal?” – but its dependence on knowing the full structural causal model is a severe bottleneck to its practical application. As such, from a causal perspective, work on RCA is broadly categorised into two groups: those that require the causal DAG, and those that do not or try to infer it from data. With the causal DAG, a widely used class of methods is traversal-based, as in Chen et al. (2014); Lin et al. (2018); Liu et al. (2021); Pham et al. (2024), wherein a variable is identified as a root cause if it is anomalous, none of its parents are anomalous, and it is linked to the target variable through a path of anomalous variables. For example, Hardt et al. Hardt et al. (2024) apply such a traversal algorithm to data from cloud computing to identify the root cause of performance drops in a microservice-based application Ma et al. (2020); Gan et al. (2021); Ikram (2023). When the causal DAG is unknown, the main idea behind the second line of work is that the anomalies themselves can support causal discovery. For example, Ikram et al. Ikram et al. (2022) treat RCA as a causal discovery problem where the goal is to find the target(s) of a soft intervention Jaber et al. (2020). Likewise, Gnecco et al. (2021); Améndola et al. (2021) describe how to use heavy-tailed distributions to infer the causal DAG for linear models. In a sense, this also amounts to using anomalies for causal discovery – if one counts points in the tail as anomalies.
1.1 Our contributions
Most methods assume one knows or can estimate the post-intervention, “anomalous”, distribution, but in real-world scenarios, it is not uncommon to have only a single sample from the distribution. In business applications, faults must be resolved rapidly, leaving no time to collect further samples. In personalised medicine, the goal is to treat each individual patient according to the unique root causes of their disease, and so while one has many samples of the healthy population, one assumes a priori that each diseased patient is unique. Even those methods that do not assume access to more than one sample from the anomalous distribution are often statistically ill-posed due to the need to estimate causal conditional probabilities or else to probe regression models in regions with low probability density. In this paper, we argue that these challenges can be overcome when the task is to identify a single root cause, and propose two simple and efficient methods for doing so. In section 2, we give the preliminaries to causal RCA and introduce ‘information theoretic’ (IT) anomaly scores upon which our methods are based. In section 3 we show that marginal anomaly scores are already sufficient for drawing conclusions about the causal relationships between anomalies: first for cause-effect pairs and then more generally when the causal DAG is a polytree. We show that marginal anomaly score ‘jumps’ along causal paths can enable identification of the root cause and propose SMOOTH TRAVERSAL, for when one knows the causal graph. We conclude Section 3 by arguing that the heuristic of selecting the variables with the highest IT anomaly scores as candidate root causes is causally justified in polytrees, and propose SCORE ORDERING, for when one does not know the causal graph. We also discuss its theoretical limitations. Finally, in Section 4 we compare existing approaches to RCA to our own proposals on both synthetic and real-world data, demonstrating that despite their simplicity, our approaches are competitive. To our knowledge, no other methods operate with causal guarantees in an RCA setting as general as we address in this paper.
2 Preliminaries
We consider the scenario in which the value $x_{n}$ of a target variable $X_{n}$ has been flagged as an anomaly by an anomaly detection algorithm. We have jointly observed values $(x_{1},\dots,x_{n})$ of variables $(X_{1},\dots,X_{n})$. Our goal is to identify a unique root cause of the anomaly $x_{n}$ from among the variables $X_{1},\dots,X_{n}$.
We assume the causal relationships between $X_{1},\dots,X_{n}$ are given by a Causal Bayesian Network (CBN) Pearl (2000), so that their joint distribution factorises into causal conditionals according to an underlying causal DAG:
$P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{i}\mid PA_{i}),$ (1)
where each $P(X_{i}\mid PA_{i})$ corresponds to the causal mechanism of variable $X_{i}$ given its parents $PA_{i}$ in the causal DAG. We then model the anomaly $x_{n}$ as having arisen due to one of the causal mechanisms having been ‘corrupted’. We observe only the single sample $(x_{1},\dots,x_{n})$ drawn from the ‘anomalous’ distribution
$\tilde{P}(X_{1},\dots,X_{n})=\tilde{P}(X_{j}\mid PA_{j})\prod_{i\neq j}^{n}P(X_{i}\mid PA_{i}),$ (2)
wherein the ‘normal’ mechanism $P(X_{j}\mid PA_{j})$ has been replaced by a corrupted one $\tilde{P}(X_{j}\mid PA_{j})$ at the root cause. Identifying the root cause is then a matter of identifying which causal conditional changed from the normal period based on a single sample from the anomalous one. While we can never exclude the case where several mechanisms are corrupted, we always start with the working hypothesis that there has been only one. If this does not work, we can still try hypotheses with more than one, but with the strong inductive bias of preferring explanations where most mechanisms worked as expected, in agreement with the so-called ‘sparse mechanism shift hypothesis’ Peters et al. (2016); Schölkopf et al. (2021). It is also important to note that although we have adopted the language of CBNs, our results can be recast in terms of the counterfactual contribution analysis introduced in Budhathoki et al. (2022) (see Appendix A). Indeed, it is noteworthy in itself that the counterfactual description can be simplified to an interventional one for the case of identifying a single root cause.
2.1 Anomaly score
We next define an anomaly score, which we use in the following sections. Let $\tau:\mathcal{X}\rightarrow\mathbb{R}$ be an appropriate feature map, which can be any existing anomaly score function, mapping elements of $\mathcal{X}$ to real values, for example, the z-score. Further, define the event
$E:=\{\tau(X_{n})\geq\tau(x_{n})\},$ (3)
of being at least as extreme as the observed event $x_{n}$ according to the feature function $\tau$. As the goal of RCA is to identify which causal mechanism has been corrupted, an important interpretation of $\tau$ is as a test statistic for the null hypothesis that $x_{n}$ was drawn from $P(X_{n})$, i.e., $H_{0}:X_{n}\sim P(X_{n})$, with $P(\tau(X_{n})\geq\tau(x_{n}))=P(E)$ as the corresponding $p$-value. Small $P(E)$ indicates that the null hypothesis is rejected with high confidence [^1] This interpretation is certainly invalid in a standard anomaly detection scenario where Eq. 3 is repeatedly computed for every observation in a sequence. However, so-called anytime $p$-values Maharaj et al. (2023) are outside of the scope of this paper. and we assume rather that the usual mechanism generating samples from $P(X_{n})$ has been corrupted for that specific statistical unit. Since small $p$-values correspond to strong anomalies, we define the marginal anomaly score by
$S(x_{n}):=-\log P(\tau(X_{n})\geq\tau(x_{n}))=-\log P(E).$ (4)
Since the logarithm of the probability of an event measures its surprise in information theory Hawkins (1980); Aggarwal (2017), as in Budhathoki et al. (2022), we call anomaly scores with the above calibration IT scores for information theoretic. Given $k$ observations $x_{n}^{1},\dots,x^{k}_{n}$, including the single anomalous one, we will use the following simple estimator [^2] This estimator is not just a finite sample approximation of (4), but also the negative logarithm of a conformal $p$-value Bates et al. (2021) for the null hypothesis of exchangeability.
$\hat{S}(x^{j}_{n}):=-\log\frac{1}{k}\left|\{i\in\{1,\dots,k\}\,:\,\hbox{with }\tau(x^{i}_{n})\geq\tau(x^{j}_{n})\}\right|.$ (5)
It is important to note that this value can be at most $\log k$, i.e., the estimated score can attain the value $s$ only when at least $e^{s}$ samples have been used. We provide detailed sample complexity results for IT anomaly score estimation in Appendix B.
In addition to the marginal anomaly score, we also introduce the conditional IT anomaly score (as in the arXiv version of Budhathoki et al. (2022)):
$S(x_{n}\mid pa_{n}):=-\log P(\tau(X_{i})\geq\tau(x_{n})\mid PA_{n}=pa_{n}),$ (6)
where now $\tau$ becomes a test statistic for the null hypothesis that $x_{n}$ was drawn according to its causal mechanism $P(X_{n}\mid PA_{n}=pa_{n})$ for that particular statistical unit. Although suppressed notationally, we allow variable-dependent feature functions $\tau_{i}$, but they must coincide between marginal and conditional anomaly scores. As the goal of RCA, as stated, is to identify which causal mechanism was corrupted, the conditional anomaly score is instrumental to what follows. To illustrate the difference between marginal and conditional anomaly scores, consider the following example:
Example 1 (Linear cause-effect model).
Consider causal model $Y:=\beta X+N$, $N\perp\!\!\!\!\perp X$. For strictly increasing $\tau$ (e.g. the z-score, $\tau(y)=(y-\mu_{Y})/\sigma_{Y}$), and observation $(x,y)$, the marginal anomaly score $S(y)=-\log P(Y\geq y)$ is simply the ‘surprise’ of the one-sided tail event for $Y$. However, the conditional anomaly score $S(y\mid x)=-\log P(Y\geq y\mid X=x)=-\log P(bx+N\geq y)=-\log P(N\geq y-bx)$, is a marginal anomaly score of the residual after accounting for $X$.
For the following sections, all proofs can be found in Appendix C unless otherwise indicated.
3 Root cause analysis when conditional probability estimation is ill-posed
To answer the question of which causal mechanism was corrupted for the sample $(x_{1},\dots,x_{n})$, we need to determine which value $x_{i}$ cannot be ‘explained’ by the values of its parents according to its normal causal conditional, $P(X_{i}\mid PA_{i})$. However, inferring the causal conditional is statistically ill-posed, particularly when the number of parents is large. One could, for example, make the additive noise assumption Hoyer et al. (2008); Peters et al. (2014) and thereby reduce the problem to that of estimating conditional expectations, but the problem remains ill-posed because analysing anomalies amounts to probing regression models in regions with low probability density. In this section, we show that causal relationships can be inferred between anomalies with a given causal DAG together with only marginal anomaly scores $\{S(x_{i}),\dots,S(x_{n})\}$, using the simple estimator in Eq. 5, in the case the causal DAG is a polytree. However, to introduce the ideas we will need, we start first with cause-effect pairs, then discuss the problem more generally in DAGs, and then lastly restrict our attention to polytrees.
3.1 Anomalies in cause and effect pairs
Let the causal DAG be $X\to Y$ for two variables $X,Y$, and we observe $(x,y)$ with anomaly scores $S(x),S(y)$. We define two null hypotheses that state that each causal mechanism, $P(X)$ or $P(Y\mid X)$, worked as normal: $H_{0}^{X}$: $x$ was drawn from $P(X)$, and $H_{0}^{Y}$: $y$ was drawn from $P(Y\mid X=x)$. Note that $H^{Y}_{0}$ does not impose a constraint on $X$, it allows $x$ to be drawn from an arbitrary distribution instead of $P(X)$, only the mechanism generating $y$ from $x$ is assumed to have remained ‘normal’. We show that whether one should reject one or both hypotheses can be determined from a comparison of $S(x)$ and $S(y)$ alone. To see how, we first need to introduce a definition which we will justify shortly:
Definition 3.1 (Bivariate score typicality).
We say that observation $(x,y)$ for $X,Y$ satisfies score typicality if
$S(y\mid x)\geq|S(y)-S(x)|_{+},$ (7)
where $|\cdot|_{+}$ denotes the positive part.
We then have the following criteria for rejecting $H_{0}^{X}$ and $H_{0}^{Y}$:
Lemma 3.2 ($p$-value bound for marginal anomaly event.).
$H^{X}_{0}$ can be rejected at level $p\leq e^{-S(x)}$.
This follows immediately from the definition of IT scores in Eq. 4. On the other hand, we obtain:
Lemma 3.3 (Anomalies rarely cause larger anomalies).
Subject to score typicality, $H^{Y}_{0}$ can be rejected at level $p\leq e^{-|S(y)-S(x))|_{+}}$.
Proof.
That $H_{0}^{Y}$ can be rejected at level $p=e^{-S(y\mid x)}$ follows from the definition of conditional IT scores in Eq. 6, the result then follows immediately, subject to score typicality. ∎
Lemma 3.3 therefore states that, subject to score typicality, an anomaly of strength $S(x)$ causes an anomaly of strength $S(y)$ only with probability at most $e^{-|S(y)-S(x)|_{+}}$, i.e. an anomaly at $X$ is unlikely to cause a much stronger anomaly at $Y$ (see also Corollary 3.3 in Ebtekar et al. (2025) for an algorithmic information theoretic view of this result).
Score typicality, therefore, gives us a criterion for rejecting $H_{0}^{Y}$ using only marginal scores, but the question remains whether we should expect it to be satisfied. The following lemma shows that for any given $y$, although we cannot guarantee score typicality is satisfied for any particular $x$, we can conclude that the bound is likely to hold approximately for a randomly chosen $x$:
Lemma 3.4 (Score typicality probably holds approximately).
Let the distribution of X (possibly multivariate) be absolutely continuous with respect to the Lebesgue measure. For any $s_{X}\leq S(y)-1$, let $x$ be chosen at random according to $P\big(X\mid S(X)\in[s_{X},S(y)]\big)$. Then we obtain, with probability at least 12/c1-2/c, an $x$ for which the following inequality holds:
$S(y\mid x)\geq|S(y)-S(x)|_{+}-2\log c.$ (8)
There are also non-trivial cases where score typicality holds exactly:
Lemma 3.5 (Injectivity and monotonicity imply score typicality).
Let the mapping $x\mapsto S(x)$ be one-to-one, and $P(S(Y)\geq S(y)\mid S(X)=S(x))\leq P(S(Y)\geq S(y)\mid S(X)\geq S(x))$. Then observation $(x,y)$ satisfies score typicality.
The first condition is the injectivity of $S$. The second is a monotonicity condition, it says: increasing the anomaly score of $X$ does not decrease the probability of an anomaly event at $Y$. Both assumptions are satisfied for all $(x,y)$, for example, for real-valued variables when $Y:=f(X,N)$, where $N$ is an independent noise term, $f$ is monotonic in $X$ and $N$ and $\tau=id$.
In the following sections, we state our results subject to score typicality for ease of exposition, but we should keep in mind that Lemmas 3.4 and 3.5 show that we could avoid it entirely, at the expense of having slightly weaker bounds holding for most samples, or by assuming injectivity and monotonicity.
Returning to our bivariate case, there is also a third hypothesis, namely that $x$ and $y$ are causally unrelated. Even if there is a causal connection between $X$ and $Y$, in this case $X\to Y$, it could happen that $x$ being anomalous was irrelevant for the anomaly $y$ because $P(Y|X=x)=P(Y)$. In other words, although $x$ is anomalous, it happens to be an anomaly that does not render an anomaly at $Y$ more likely. The corresponding null hypothesis is $H^{XY}_{0}$: $x$ was drawn from $P(X)$ and $y$ from $P(Y)$. We find:
Lemma 3.6 (Bound on $p$-value for independent anomalies).
$H^{XY}_{0}$ can be rejected at level $p\leq e^{-S(x)-S(y)}[1+S(x)+S(y)]$.
The lemma follows from the fact that for independent variables $X,Y$, the corrected sum $S(x)+S(y)-\log[1+S(x)+S(y)]=:\tilde{S}((x,y))$ is a valid IT score for pairs and thus satisfies $P(\tilde{S}((X,Y))\geq\tilde{S}((x,y)))\leq e^{-\tilde{S}((x,y))}$ (see Theorem 1 in the arXiv version of Budhathoki et al. (2022)).
Lemmas 3.2 to 3.6 illustrate some of the ideas that will be of use in the case the causal graph is a polytree: in particular, Lemma 3.3 is noteworthy as it shows that we can test hypotheses concerning conditional distributions using differences between marginal anomaly scores, so long as we have the causal graph. However, they also entail interesting consequences for the case where the causal direction is not known. For example, let $S(x)=10,S(y)=5$ and significance level $\alpha=0.01$. For $X\to Y$ we can only reject $H_{0}^{X}$, while the mechanism $P(Y|X)$ possibly worked as expected. However, for $Y\to X$, we would reject both hypotheses, that $P(Y)$ and that $P(X|Y)$ worked as expected with $p$-value $e^{-5}$ each. Following the working hypothesis that at most one mechanism was corrupted, we thus prefer $X\to Y$. Note that the third alternative of causally independent generation of $x$ and $y$ can be rejected even at level $p\leq e^{-5-10}(1+5+10)$.
3.2 Anomalies in DAGs
Before restricting our attention to the case where the causal DAG is a polytree, let us first consider the problem in full generality in DAGs. To infer for which variable the causal mechanism has been corrupted, we start with the hypothesis that all mechanisms worked as expected except for $P(X_{j}\mid PA_{j})$: $H^{j}_{0}$: for the anomaly event $(x_{1},\dots,x_{n})$ all $x_{i}$ with $i\neq j$ have been drawn from $P(X_{i}\mid PA_{i}=pa_{i})$. We will return to marginal anomaly scores shortly, but first, to define test statistics, we focus on conditional anomaly scores. We also need to make the assumption:
  1. 1.
    Continuity: all variables $X_{i}$ are continuous with density w.r.t. the Lebesgue measure, and also all conditional distributions $P(X_{i}|PA_{i}=pa_{i})$ have such a density.
Referring to Eq. 6, conditional scores with variable inputs $X_{i},PA_{i}$ define random variables $S(X_{i}\mid PA_{i})$. Continuity ensures that all conditional anomaly scores are distributed according to the density $p(s)=e^{-s}$ for $s\geq 0$, and entails a property that will be convenient for testing $H^{j}_{0}$:
Lemma 3.7 (Conditional anomaly scores are independent).
Subject to the continuity assumption, $\left\{S(X_{1}|PA_{1}),\cdots,S(X_{n}|PA_{n})\right\}$ are independent random variables.
Independence of conditional scores enables the definition of a simple test statistic that is obtained by summing over them [^3] Note that this corresponds to Fisher’s method Fisher (1925) for aggregating $p$-values, which also considers the sum of independent, log-transformed $p$-values. together with a correction term. With $S_{\rm sum}:=\sum_{i\neq j}S(X_{i}|PA_{i})$, we define the joint anomaly score
$S:=S_{\rm sum}-\log\sum_{l=0}^{n-2}\frac{S_{\rm sum}^{l}}{l!}.$ (9)
To understand Eq. 9, note that $S_{\rm sum}$ no longer has the properties of an IT anomaly score: the sum of independent IT scores is likely to be large because it is not unlikely that the set contains at least one large term. The second term is therefore needed to ‘recalibrate’ the sum and is akin to a multiple testing correction. The following result states that this is quantitatively the right correction:
Lemma 3.8 (The joint score is an IT anomaly score).
If $H^{j}_{0}$ holds, Eq. 9 is distributed according to the density $p(s)=e^{-s}$ for $s\geq 0$.
As a direct result of the above lemmas, we have:
Theorem 3.9 ($p$-value for joint anomaly event).
$H^{j}_{0}$ can be rejected for the observation $(x_{1},\dots,x_{n})$ with $p$-value $p=e^{-S_{\rm sum}}\cdot\sum_{l=0}^{n-2}\frac{S_{\rm sum}^{l}}{l!}$.
The theorem justifies choosing the index $j$ with the maximal conditional anomaly score as the root cause, whenever we follow the working hypothesis that only one mechanism is corrupted.
With our framework in place, we can return to our original goal: to show how RCA is possible, without the estimation of conditional probabilities, using only marginal anomaly scores. In order to do so, we restrict our attention to the case where the causal DAG is a polytree.
3.3 Anomalies in polytrees
Polytrees are DAGs where the underlying skeleton is a tree (see Fig. 2 in Appendix D). The key property of polytrees that we exploit is that the parents of any given variable are marginally independent. This allows us to replace conditional anomaly scores with bounds derived from marginal anomaly scores, following the ideas from the bivariate case. To this end, note that, as they are independent, we may use the same construction as we did in Eq. 9 to define a joint score over the parents of a given variable. For variable $X_{i}$, with parents $PA_{i}^{1},\dots,PA_{i}^{k}$, and event $pa_{i}$:
$S_{\rm joint}(pa_{i}):=\sum_{j=1}^{k}S(pa_{i}^{j})-\log\sum_{l=0}^{k-1}\frac{(\sum_{j=1}^{k}S(pa_{i}^{j}))^{l}}{l!}.$ (10)
That this is indeed an IT score for joint event $pa_{i}$ follows from the same proof as for Lemma 3.8. For brevity, we drop the subscript and simply write $S(pa_{i})$ to mean the joint anomaly score of multivariate event $pa_{i}$, as it will be clear from context whether we refer to a marginal score or a joint score. By using the joint score over the parents of each variable, we have effectively reduced the polytree case to the bivariate one with respect to each variable and its parents. In particular, we generalise score typicality:
Definition 3.10 (Score typicality).
We say that observation $(x,pa_{i})$ for $X,PA_{i}$ satisfies score typicality if
$S(x_{i}\mid pa_{i})\geq|S(x_{i})-S(pa_{i})|_{+}.$ (11)
Subject to the continuity assumption, Lemma 3.4 applies directly, in statement and proof, to score typicality as defined above, simply substituting $Y$ for $X_{i}$ and $X$ for $PA_{i}$. We can therefore say that for observation $x_{i}$, score typicality is likely to hold approximately for a randomly chosen $pa_{i}$, in the sense given by Lemma 3.4.
As a result, we obtain the following bound on the joint statistics introduced in the previous subsection:
Theorem 3.11.
(Bound on $p$-value for joint anomaly event using only marginal anomaly scores). With $\hat{S}_{\rm sum}=\sum_{i\neq j}|S(x_{i})-S(pa_{i})|_{+}$ it holds that subject to score typicality, $H^{j}_{0}$ can be rejected with $p$-value
$p\leq e^{-\hat{S}_{\rm sum}}\cdot\sum_{l=0}^{n-2}\frac{\hat{S}_{\rm sum}^{l}}{l!}.$ (12)
We conclude that $H^{j}_{0}$ needs to be rejected whenever the anomaly score increases significantly along a path of anomalies at any $i\neq j$. This justifies inferring the index $j$ that maximises the score difference $|S(x_{j})-S(pa_{j})|_{+}$ as the unique root cause (with the difference being $|S(x_{i})-0|_{+}=S(x_{i})$ for any node with no parents) because this yields the weakest bound in Eq. 12. Motivated by these bounds, we propose the algorithm SMOOTH TRAVERSAL [^4] We thank Elke Kirschbaum for suggesting this name. which selects the node as the root cause that shows the strongest increase in anomaly score compared to its parents (as a heuristic we consider the score difference with the most anomalous parent of each node, see Algorithm 1 below). In the usual traversal approach Chen et al. (2014); Lin et al. (2018); Liu et al. (2021), a variable is identified as a root cause if it is anomalous, none of its parents are, and it is linked to the target variable through a path of anomalous variables. However, this requires the user to choose a threshold above which a node is considered anomalous. SMOOTH TRAVERSAL avoids having to make any such arbitrary choice. We can prove that if the maximum score difference is large, then the node at which the difference is maximised is likely to be the root cause. To this end: for anomaly event $(x_{1},\dots,x_{n})$ and anomaly score differences
$\delta_{i}:=\big|S(x_{i})-S(pa_{i})\big|_{+},\quad i=1,\dots,n,$
define null hypothesis $H_{0}^{\max}$: for anomaly event $(x_{1},\dots,x_{n})$ and $j={\rm argmax}_{i}\,\delta_{i}$, $x_{j}$ was drawn from $P(X_{j}\mid PA_{j}=pa_{j})$, i.e. the anomaly with the largest score difference with its parents is not the root cause.
Theorem 3.12 (Bound on $p$-value for anomaly with biggest jump not being the root cause).
Subject to score typicality, when there is a single root cause, $H_{0}^{\rm max}$ can be rejected for observation $(x_{1},\dots,x_{n})$, and maximum score difference $\delta_{\max}:={\rm max}(\delta_{1},\dots,\delta_{n})$ with $p$-value
$p\leq 1-(1-e^{-\delta_{\rm max}})^{n-1}.$ (13)
The theorem follows from the fact that if the node maximising the score difference from its parent is not the root cause, it implies that at least one conditional anomaly score, of the (n1)(n-1) non-root causes, is at least $\delta_{\rm max}$. [^5] The approach for deriving this $p$-value broadly resembles Tippett’s method Tippett (1931) for combining independent $p$-values.
Algorithm 1 (Smooth Traversal) Returns the variable with the largest positive score difference to its highest scoring parent
0: Scores $S(x_{1}),\ldots,S(x_{n})$, causal DAG $G$ of variables $X_{1},\dots,X_{n}$
0: Variable $X_{k}$ with the largest positive score difference to its highest-scoring parent
1: Initialize $\text{max\_min\_difference}\leftarrow$-$\infty$
2: Initialize $\text{candidate\_root\_cause}\leftarrow\text{None}$
3:for each ancestor $X_{i}$ of $X_{n}$ in $G$ do
4:  Let $\text{PA}_{i}$ be the set of parents of $X_{i}$ in $G$
5:  if $\text{PA}_{i}\neq\emptyset$ then
6:   $\text{max\_parent\_score}\leftarrow\max\limits_{X_{j}\in\text{PA}_{i}}S(x_{j})$
7:  else
8:   $\text{max\_parent\_score}\leftarrow 0$
9:  end if
10:  $\text{score\_difference}\leftarrow|S(x_{i})-\text{max\_parent\_score}|_{+}$
11:  if $\text{score\_difference}>\text{max\_min\_difference}$ then
12:   $\text{max\_min\_difference}\leftarrow\text{score\_difference}$
13:   $\text{candidate\_root\_cause}\leftarrow X_{i}$
14:  end if
15:end for
16:return candidate_root_cause
3.4 Shortlist of root causes via ordering anomaly scores
We have shown that a unique root cause can be identified using only marginal anomaly scores if the causal graph is known. We now drop the assumption that the causal graph is known and ask, can we find the root cause given only the marginal anomaly scores $S(x_{1}),\dots,S(x_{n})$? We argue that the top-scoring anomalies constitute a good shortlist for the root cause.
To see this, assume again that the (unknown) causal DAG for variables $X_{1},\dots,X_{n}$ is a polytree as before, and for illustration, first consider a particular causal path through the graph,
$X_{i_{1}}\to X_{i_{2}}\to\dots\to X_{i_{k}}.$
If we now ‘coarsen’ the graph, marginalising out all the intermediate variables between $X_{i_{1}}$ and $X_{i_{k}}$, so that we collapse the whole path into a single edge $X_{i_{1}}\to X_{i_{k}}$, application of Lemma 3.3 shows that we can reject the hypothesis that the new ‘coarse’ mechanism $P(X_{i_{k}}\mid X_{i_{1}}=x_{i_{k}})$ worked as expected at level $p\leq e^{-|S(x_{i_{k}})-S(x_{i_{1}})|_{+}}$. However, note that the coarse mechanism for $X_{i_{k}}$ now implicitly encompasses each of the intermediate mechanisms that connected $X_{i_{1}}$ to $X_{i_{k}}$ in the full graph. In other words, the anomaly $x_{i_{1}}$ is unlikely to cause a much larger anomaly $x_{i_{k}}$ anywhere downstream, unless we allow that any of the intermediate mechanisms connecting them has been corrupted.
With this in mind, consider the full graph on $X_{1},\dots,X_{n}$ and let $\pi$ be a permutation of $\{1,\dots,n\}$ such that
$S(x_{\pi(1)})\geq S(x_{\pi(2)})\geq\dots\geq S(x_{\pi(n)}).$
We define the null hypothesis that the mechanisms worked as expected for all of the top-$k$ scoring anomalies: $H_{0}^{\text{top-}k}$: for anomaly event $(x_{1},\dots,x_{n})$ all $x_{\pi(i)}$ with $i\leq k$ have been drawn from $P(X_{\pi(i)}\mid PA_{\pi(i)}=pa_{\pi(i)})$. Under the null hypothesis, we know that the score for $x_{\pi(1)}$ must have been greater by at least $\Delta_{k}:=S(x_{\pi(1)})-S(x_{\pi(k+1)})$ compared to an ancestral node, despite its mechanism, and the mechanisms of any intermediates, working as expected. This is because whether or not any of the variables $X_{\pi(2)},\dots,X_{\pi(k)}$ are intermediates, their mechanisms, by assumption, worked as expected under the null. If any of the variables $X_{\pi(k+2)},\dots,X_{\pi(n)}$ are intermediates (or if $X_{\pi(1)}$ has no ancestors), then its score increased by an even greater amount than $\Delta_{k}$. Applying the same coarsening idea as above, and accounting for multiple testing, we therefore have:
Theorem 3.13.
(Bound on likelihood root cause not in top-$k$ highest scoring anomalies). Subject to score typicality, where $d_{\rm max}$ denotes the maximum in-degree among all nodes in the graph, we can reject $H_{0}^{\textnormal{top-}k}$ at level $p\leq n\cdot d_{\rm max}\cdot e^{-\Delta_{k}}$.
These bounds motivate SCORE ORDERING which, given the marginal anomaly scores and an assumed upper bound on the in-degree of nodes in the causal graph, returns the set of top-$k$ highest scoring anomalies with the guarantee that the root cause is among them with at least the desired confidence level $1-\alpha$ (see Algorithm 2). Note that one could just as well specify a desired number of variables, $k$, and SCORE ORDERING would then return a confidence level for each possible $d_{\rm max}$ so that the latter is not a strict input requirement.
Algorithm 2 (Score Ordering) Returns the top-$k$ scoring anomalies such that the root cause is among them with confidence at least $1-\alpha$
0: Scores $S(x_{1}),\ldots,S(x_{n})$, maximum in-degree $d_{\rm max}$, confidence level $1-\alpha$
0: A set $L\subseteq\{X_{1},\ldots,X_{n}\}$
1: Let $\pi$ be a permutation of $\{1,\ldots,n\}$ such that $S(x_{\pi(1)})\geq S(x_{\pi(2)})\geq\cdots\geq S(x_{\pi(n)})$
2: Initialize $L\leftarrow\emptyset$
3:for $k=1$ to $n$ do
4:  $L\leftarrow L\cup\{X_{\pi(k)}\}$
5:  if $k=n$ then
6:   return $L$
7:  end if
8:  $\Delta\leftarrow S(x_{\pi(1)})-S(x_{\pi(k+1)})$
9:  if $n\cdot d_{\rm max}\cdot e^{-\Delta}\leq\alpha$ then
10:   return $L$
11:  end if
12:end for
3.5 Beyond polytrees: limitations of score ordering
Most of the theory developed in this section has depended on the assumption that the causal graph is a polytree. While this assumption appears in causal literature (e.g.  Tramontano et al. (2022); Jakobsen et al. (2022); Choo et al. (2024)), it is nonetheless a strong structural assumption. One of the strengths of both of the algorithms we propose, SMOOTH TRAVERSAL and SCORE ORDERING, is their simplicity and efficiency. SCORE ORDERING in particular demands very weak input requirements, and so the fact that it comes with any causal guarantees at all might be a surprise. As a minimum, this might justify its use as a ‘first pass’ heuristic in settings beyond those we study here. However, the question remains as to what extent we might expect our proposed algorithms to perform well in settings that do not meet the structural assumptions (we empirically assess this question in the next section). The key observation on which SCORE ORDERING depends in the proof of Theorem 3.13 is that an anomaly $x_{i}$ is unlikely to cause a much larger anomaly downstream. Li et al. Li et al. (2024) describe a condition under which the $z^{2}$-score of an effect $X_{j}$ can be larger than the $z^{2}$-score of the root cause $X_{i}$ in a DAG with linear structural equations. Although our results refer to IT scores rather than $z^{2}$-scores, their example nonetheless applies to us as the $z^{2}$-score can be turned into an IT score by a monotonic recalibration. They show (see Corollary 2.1 in Li et al. (2024)) that a sufficient condition for such a case not to occur is that the DAG be a polytree, and so their results do not contradict our own. However, in addition, we show in Appendix E that in DAGs (with linear structural equations) more generally such cases are ‘rare’ in the sense that, under reasonable assumptions, it will hold for only a ‘small’ fraction of root causes. Despite, then, exceptions occurring, we believe our results nonetheless give confidence that SCORE ORDERING can be applied as a useful first heuristic even beyond polytrees.
4 Experiments
We evaluate SMOOTH TRAVERSAL and SCORE ORDERING against a range of existing RCA methods, that can also operate with a single anomalous sample, on both synthetic and real world data. In particular we evaluate our methods against:
  1. ‘Cholesky’ Li et al. (2024) Assumes linear structural causal models, but does not require the causal graph. Makes use of the Cholesky decomposition of the covariance matrix of observed variables, exploiting an invariance property across permutations to identify the root cause.
  2. ‘Traversal’ Liu et al. (2021); Chen et al. (2014); Lin et al. (2018) Requires the causal graph. Identifies nodes which are anomalous, have no anomalous parents, and are linked to the target node via a path of anomalous nodes as root causes.
  3. Circa Li et al. (2022) Requires the causal graph. Fits a linear structural causal model to the data from the normal period and returns the node with the largest residual, given its parents, in the anomalous period.
  4. ‘Counterfactual’ Budhathoki et al. (2022) Requires the structural causal model. Finds the counterfactual Shapley contribution of each node to the anomalousness of the target event and outputs the node with the highest contribution.
We also evaluate against RCD Ikram et al. (2022) and $\varepsilon$-Diagnosis Shan et al. (2019) in Appendix F, however, neither method operates in the single anomalous sample setting, and so we do not recommend direct comparison.
Synthetic data generation: full details are provided in Appendix F. In brief, to generate data, we sample a random DAG (not restricted to a polytree) of $n$ nodes. Structural equations are randomly assigned as simple feed-forward neural networks or as linear models of the parents and noise variables. In all experiments, 1000 observations are drawn according to the randomly assigned model in the normal period. To produce an anomalous sample, a root cause is chosen at random from among the nodes and a target node from among its descendants (including itself). An anomaly is then injected at the root cause by adding $x$ multiples of its standard deviation to its value, and propagated through the causal model. In each experiment, this process is repeated 100 times per data point. Unless otherwise stated, each algorithm is evaluated by its top-1 recall. With synthetic data, we answer the following questions:
How does performance vary with anomaly strength? We display results for $x\in\{2.0,2.1,2.2,\dots,3.0\}$ in Fig. 1. We find that the strongest performing algorithms are SMOOTH TRAVERSAL, Traversal, and Counterfactual, all of which outperform SCORE ORDERING. Circa and Cholesky perform considerably worse than the other approaches, apparently due to the assumption of linearity (see Fig. 5 in Appendix F). The performance of SCORE ORDERING improves slightly as the anomaly strength increases, but performance is approximately constant for the other approaches. We extend this analysis in Appendix F.2: to RCA and $\varepsilon$-Diagnosis (Fig. 4), using only linear SCMs (Fig. 5), and restricting the DAG to being a polytree (Fig. 6).
Figure 1: True positive rate for identifying the root cause against anomaly strength injected at the root cause.
How do runtimes scale with increasing graph size? For a fixed $x=3$, we measure the runtime of each algorithm for graphs with an increasing number of nodes, generated as described above. We observe that (see Fig. 3 in Appendix F), Traversal and SMOOTH TRAVERSAL are the fastest, with the remaining approaches having comparable average runtimes.
How does performance vary with increasing graph size? We generate causal models with $n\in\{20,40,\ldots,100\}$ and fixed anomaly strength of $x=3$ (see Fig. 7). While the performance of SMOOTH TRAVERSAL, Traversal and Counterfactual seems to remain constant, the performances of SCORE ORDERING, Cholesky, and Circa all begin to decrease as the graph increases in size.
How robust are the methods to misspecification of the causal graph? For the approaches that require a causal graph, we investigated how their performances varied as an increasing number of random edge additions/removals/reversals were introduced to the given graph (see Appendix F.4). For $n=50$ and $x=3.0$, we see SMOOTH TRAVERSAL, Traversal and Counterfactual are similarly robust, showing similar delays in performance, while the performance of Circa decayed more rapidly.
Real world data evaluation: We additionally performed a comprehensive evaluation with two ‘real-world’ datasets: the PetShop Hardt et al. (2024) dataset (see Appendix F.6) and the Sock-shop 2 dataset Pham et al. (2024) (see Appendix F.7), as well as ‘semi-synthetic’ datasets generated using the ProRCA package Dawoud and Talupula (2025) (see Appendix F.8). In these evaluations, the picture can be quite different from that in the synthetic data experiments. SCORE ORDERING often performs well across the datasets, but both of our algorithms are variously outperformed by Cholesky, Circa or Counterfactual in certain settings.
5 Conclusion
We have explored several directions in which the practical limitations of RCA can be addressed: first by avoiding estimation of conditional distributions in the setting that the causal graph is known, and second by avoiding needing the causal graph at all, in the setting that it is a polytree. To do so, we leverage information-theoretic (IT) anomaly scores and prove general laws about their typical decay along causal paths without any parametric assumptions. Using our results, we then propose two novel algorithms: SMOOTH TRAVERSAL and SCORE ORDERING, give probabilistic guarantees for the correctness of their outputs, and show they have competitive performance in both synthetic and real-world data despite their simplicity. To our knowledge, our work is the only which provides non-parametric guarantees for RCA with a single anomalous point without requiring the full structural causal model.
Acknowledgments and Disclosure of Funding
We would like to thank Manuel Gomez-Rodriguez for fruitful discussions during different stages of this project. W. R. Orchard would like to thank Peterhouse, Cambridge, and Cancer Research UK for their support for this work.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All the theoretical claims in the abstract and introduction are part of the main article.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Most of the limitations come from theoretical assumptions made to obtain the results. These limitations are discussed in the main article.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Most of the article is given over to theoretical results and their proofs. We clearly state all assumptions, and proofs of every lemma and theorem in the paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We give a detailed description of the set-up for all experiments performed, and point to 3rd party Python libraries where applicable.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide a link to the complete code for reproducing our experimental results at the end of the article.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, as in answer to questions 4 and 5 above.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In all figures where error bars are necessary and large enough to be visible, they are displayed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The work requires very modest compute resources, but details have been provided in the appendix nonetheless.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Due to the paper being entirely theoretical, the paper adheres to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There are no broad societal impacts of the application of the ideas outlined in this paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No major risks are posed by the paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The only assets used by the paper are public datasets. We cite all original papers and a URL to access it alonside licensing information.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The only assets are code for running the experiments (to be released on acceptance at the latest). This code is well documented. The paper otherwise released no new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The paper does not involve the use nor development of LLMs in any way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Relationship to contribution analysis
We first sketch the causal framework adopted by Budhathoki et al. Budhathoki et al. (2022). In particular, they assume that the causal relationships between $X_{1},\dots,X_{n}$ are described by a Structural Causal Model (SCM). In an SCM, each variable $X_{i}$ is a function $f_{i}$ of its parents $PA_{i}$ in the causal graph, and an unobserved noise term $N_{i}$:
$X_{i}:=f_{i}(PA_{i},N_{i}),$ (14)
where the noise variables $N_{1},\dots,N_{n}$ are jointly independent Pearl (2000). In addition, it is assumed that the SCM is invertible Zhang et al. (2015). That is, the noise value $n_{j}$ of $N_{j}$ can be recovered from the value $x_{j}$ of its corresponding observed variable $X_{j}$ and the values $pa_{j}$ of its parents.
As in our case, the value $x_{n}$ is flagged as an anomaly, and we wish to identify its root cause from among the variables $(X_{1},\dots,X_{n})$. Treating (without loss of generality) $X_{n}$ as a sink node in the causal DAG, iterative application of Eq. 14 results in the representation
$X_{n}=F(N_{1},\dots,N_{n}),$ (15)
in which the structural information is implicit in the function $F$. By Eq. 15 we therefore have that $x_{n}=F(n_{1},\dots,n_{n})$ where $(n_{1},\dots,n_{n}):=\mathbf{n}$ denote the corresponding values of the noise variables for sample $(x_{1},\dots,x_{n})$. This representation makes clear that a node $X_{j}$’s contribution to the anomaly $x_{n}$ is ultimately attributable to the contribution of its noise term Budhathoki et al. (2022); Von Kügelgen et al. (2023).
If you consider our presentation of RCA, wherein we attribute the anomaly $x_{n}$ to one of the causal mechanisms being corrupted, the characterisation in terms of noise terms may appear quite different. However, we can connect the two perspectives by noting that one can think of each value $n_{j}$ as randomly switching between deterministic mechanisms $f_{j}(\cdot,n_{j})$ – so-called response functions Balke and Pearl (1994). If a noise term is "corrupted" and takes an unusual value $n_{j}$, the corresponding mechanism generating $x_{j}$ from its parents $pa_{j}$ will be unusual as well. The goal of finding which causal mechanism did not work as expected, then is common between the two perspectives. The intuition behind the approach of Budhathoki et al. is that if for the root cause $X_{j}$ the unusual value of its noise term $n_{j}$ were replaced by a "normal" one, then $x_{n}$ would change to a non-anomaly with high probability. This gives the basis for defining how much each variable contributed to the observed anomaly. We explain how this is formalised next.
A.1 Contribution analysis
To compute the contribution of each noise $N_{i}$ to the anomaly $x_{n}$, Budhathoki et al. Budhathoki et al. (2022) measure how replacing the observed value $n_{i}$ (originating from a potentially corrupted mechanism) with a random "normal" value, sampled from its regular distribution $P(N_{i})$, changes the likelihood of the anomaly event, $E=\{\tau(X_{n})\geq\tau(x_{n})\}$. Intuitively, this shows us the extent to which $n_{i}$ was responsible for the extremeness of $x_{n}$.
Note, however, that the influence of the value $n_{i}$ on the likelihood of the anomaly event will also depend on the values of the other noise variables. In order to assess the contribution of $n_{i}$ we therefore also need to consider how it depends on the context, i.e. how it changes depending on which other noise variables we similarly replace with random "normal" values. To formalise this, for any subset of the index set, ${\mathcal{S}}\subseteq\mathcal{U}:=\left\{1,\ldots,n\right\}$, first note that the probability of the anomaly event when all nodes in $\mathcal{S}$} are set to their observed value and all the nodes in $\bar{{\mathcal{S}}}=\mathcal{U}\setminus{\mathcal{S}}$ are randomised is $P(E\,|\,{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}})$. Now the contribution of a node $j\not\in{\mathcal{S}}$ given the context $\mathcal{S}$} is defined as: [^6] Note the difference in the definition of $C(.|{\mathcal{S}})$ compared to Budhathoki et al. (2022). In our case $\mathcal{S}$} is the set for which ${\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}}$ while in Budhathoki et al. (2022) ${\bf N}_{\mathcal{U}\setminus{\mathcal{S}}}={\bf n}_{\mathcal{U}\setminus{\mathcal{S}}}$.
$C(j|{\mathcal{S}}):=\log\frac{P(E|{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}},{\bf N}_{j}={\bf n}_{j})}{P(E|{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}})}.$ (16)
To give a "fair" attribution among the noise variables, the authors adopt Shapley values. Let $\Pi:\mathcal{U}\rightarrow\mathcal{U}$ be the set of all possible permutations of the nodes and $\pi\in\Pi$ be any permutation. One then defines the contribution of a node $j$ given permutation $\pi$ as $C^{\pi}(j):=C(j|I^{\pi<j})$, where $I^{\pi<j}$ denotes the set of nodes that appear before $j$ with respect to $\pi$, i.e., $I^{\pi<j}=\left\{i\in\mathcal{U}\,|\,\pi(i)<\pi(j)\right\}$. We easily see that for each permutation $\pi$, $S(x_{n})$ decomposes into the contributions of each node, i.e., $S(x_{n})=\sum_{j\in\left\{1,\ldots,n\right\}}C^{\pi}(j)$. The Shapley contribution of a node is then given by averaging over all the possible permutations in $\Pi$:
$C^{Sh}(j):=\frac{1}{n!}\sum_{\pi\in\Pi}C(j|I^{\pi<j}).$ (17)
This approach therefore provides a full quantitative contribution analysis of root causes. However, it suffers from some practical issues. Firstly, the Shapley contribution of a node is expensive to compute Budhathoki et al. (2022). More fundamentally, however, for most permutations $\pi$, the contributions $C^{\pi}(j)$ rely on knowing the structural equations (14). As the SCM cannot generally be inferred even with interventional data, this is a serious bottleneck for the application of this approach. As the approaches we propose in this paper do not depend on knowing the structural equations (or even the causal graph in the case of SCORE ORDERING), we explain next how we can nonetheless recast our results in terms of the contribution approach just presented.
A.2 Interventional vs counterfactual RCA
The key result we will need is that so long as $\pi$ is a topological ordering of the nodes in the causal graph, then all contributions $C^{\pi}(j)$ do not require knowing the SCM. To illustrate this, first consider the bivariate case $X\to Y$ with structural equations $X:=N_{X}$ and $Y:=f(X,N_{Y})$, and anomaly $y$ in sample $(x,y)$. To determine the contribution of $N_{Y}$ to the anomaly, we consider randomising $N_{Y}$ and fixing $N_{X}=n_{X}=x$. This generates $Y$ according to the observational distribution $P(Y\mid x)$, so its estimation does not require knowledge of the SCM. On the other hand, randomising $N_{X}$ and fixing $N_{Y}=n_{Y}$ cannot be resolved into any observational term (see also Section 5 in Janzing et al. (2024)). [^7] This can be checked by an SCM with two binaries where $Y:=X\oplus N_{Y}$, with unbiased $N_{Y}$, which induces the same distribution as the SCM $Y:=N_{Y}$, where $X$ and $Y$ are disconnected and thus $X$ has a contribution of zero.
The following result generalises this insight:
Proposition A.1.
Whenever $\pi$ is a topological order of the causal DAG, i.e., there are no arrows $X_{i}\to X_{j}$ for $\pi(i)>\pi(j)$, all contributions $C^{\pi}(j)$ can be computed from observational conditional distributions.
This will allow us to connect our approach to that of Budhathoki et al. Budhathoki et al. (2022) in the following sense: if our goal is to simply identify a single root cause, then as long as it is "dominating" in the sense that it has the largest contribution of any variable regardless of the order chosen, Proposition A.1 says we can identify it without knowing the SCM.
It will prove instructive to generalise the notion of the contribution of a single node in (16) to the contribution of a set $\mathcal{R}\subseteq\mathcal{U}\setminus{\mathcal{S}}$ of nodes, given the context $\mathcal{S}$}, i.e.,
$C(\mathcal{R}|{\mathcal{S}}):=\log\frac{P(E|{\bf N}_{\mathcal{R}}={\bf n}_{\mathcal{R}},{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}})}{P(E|{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}})}.$ (18)
This notion becomes rather intuitive after observing that it is given by the sum of the contributions of all the elements in $\mathcal{R}$ when they are one by one added to the context $\mathcal{S}$}:
Lemma A.2.
For any set $\mathcal{R}\subseteq\mathcal{U}\setminus{\mathcal{S}}$, it holds that $C(\mathcal{R}|{\mathcal{S}}):=C(j_{1}|{\mathcal{S}})+\sum_{i=2}^{k}C(j_{i}|{\mathcal{S}}\cup\left\{j_{1},\dots,j_{i-1}\right\})$ with $\mathcal{R}=\left\{j_{1},\ldots,j_{k}\right\}$.
Next, the following result shows that a set of nodes are unlikely to obtain a high contribution when the noise values are randomly drawn from their usual distributions, i.e., we have the following proposition:
Proposition A.3.
Whenever all noise variable in some set $\mathcal{R}$ are sampled from $P({\bf N}_{\mathcal{R}})$, it holds that $P\left(C(\mathcal{R}|{\mathcal{S}})\geq\alpha\right)\leq e^{-\alpha}$.
Note that the proposition allows that the variables $N_{j}$ not in $\mathcal{R}$ are drawn from a different distribution. Thus, Proposition A.3 states that it is unlikely that a set that only contains non-corrupted nodes has high contribution.
Next, we will show how our main results can be recast in terms of bounds on contributions, rather than on $p$-values.
A.3 Bounds on contributions
Consider, for illustration, the setting discussed in subsection 3.1 wherein the causal DAG is $X\to Y$, we observe $(x,y)$ with anomaly scores $S(x),S(y)$, and we wish to find the root cause of anomaly $y$. Under the same assumptions as stated there we have the following proposition:
Proposition A.4.
Subject to score typicality, the contributions of $y$ and $x$ on the anomaly $y$ satisfy:
$C(x)\leq S(x)\quad\text{and}\quad C(y\mid x)\geq|S(y)-S(x)|_{+}$
Proof.
Proposition A.1 permits rewriting contributions in terms of observational probabilities, so that, together with score typicality, we have:
$C(y\mid x):=$ $\log\frac{P(\tau(Y)\geq\tau(y)\mid Y=y,X=x)}{P(\tau(Y)\geq\tau(y)\mid X=x)}$
$=\log\frac{1}{P(\tau(Y)\geq\tau(y)\mid X=x)}$
$=S(y\mid x)\geq|S(y)-S(x)|_{+},$
and using $C(x)+C(y\mid x)=S(y)$ we obtain the bound for $C(x)$. ∎
We can now further interpret the conclusions drawn in subsection 3.1, in terms of contributions. For example, suppose we have $S(y)\gg S(x)$. In the case that $X\to Y$, Proposition A.4 says that $y$ must have a large contribution to the anomaly, while $x$ must have a smaller one. We would therefore favour $y$ as the root cause. Alternatively, if $S(x)\gg S(y)$, then we cannot conclude that $y$ has a large contribution, but we would reject the hypothesis that the mechanism generating $x$ worked as normal (from Lemma 3.2), and its contribution may be large. Under the working hypothesis that there is only a single root cause, we would favour $x$. Following the extensions from the bivariate case in subsections 3.2 and 3.3, we can similarly recast the results in terms of contributions as above. Naturally, this does not alter any of the conclusions, but rather demonstrates that our results do not depend strictly on our adopted formalisation of the RCA problem in terms of CBNs, but additionally admits an interpretation in terms of contributions.
Appendix B Sample complexity of information-theoretic anomaly score estimation
To reliably tell whether the IT anomaly score of one anomaly exceeds that of another, we need to estimate each score up to an error of at least half their difference. We investigate how many samples are required to estimate the scores $S(x_{i}),\dots,S(x_{n})$ up to an error level $\delta$:
Lemma B.1.
If we use at least
$k=\frac{3e^{S_{\max}}}{\delta^{2}}\log\left(\frac{2n}{\alpha}\right)$ (19)
samples from the ‘normal’ period, the score estimates $\hat{S}(x_{1}),\dots,\hat{S}(x_{n})$ using the estimator in Equation (5) satisfy $|\hat{S}(x_{i})-S(x_{i})|<\delta$ for all $i\in\{1,\dots,n\}$ with probability at least $1-\alpha$.
Proof.
Let $p_{i}:=P(\tau(X_{i})\geq\tau(x_{i}))$ and estimate $\hat{p}_{i}:=\frac{1}{k}\sum_{j=1}^{k}\mathbf{1}\{\tau(x_{i}^{j})\geq\tau(x_{i})\}$ so that $S(x_{i})=-\log p_{i}$ and $\hat{S}(x_{i})=-\log\hat{p}_{i}$ as in Equations (4) and (5), respectively. $|\hat{S}(x_{i})-S(x_{i})|<\delta$ is equivalent to,
$|\log\hat{p}_{i}-\log p_{i}|<\delta\Longrightarrow e^{-\delta}<\frac{\hat{p}_{i}}{p_{i}}<e^{\delta}.$ (20)
Rearranging, we have $p_{i}(e^{-\delta}-1)<\hat{p}_{i}-p_{i}<p_{i}(e^{\delta}-1)$, and noting $e^{\delta}-1>1-e^{-\delta}$ for all $\delta>0$, we have
$|\hat{p}_{i}-p_{i}|<\epsilon_{i}:=(e^{\delta}-1)p_{i},$ (21)
so that when $\delta$ is small we have $\epsilon_{i}\approx\delta p_{i}$. We can then use the multiplicative form of the Chernoff bound for $0\leq\delta\leq 1$ to say,
$P(|\hat{p}_{i}-p_{i}|\geq\delta p_{i})\leq 2\exp\left(-\frac{\delta^{2}p_{i}k}{3}\right).$ (22)
Noting that $\min_{i}p_{i}=\min_{i}e^{-S(x_{i})}=e^{-\max_{i}S(x_{i})}=:e^{-S_{\rm max}}$ and applying the union bound,
$P\left(\bigcup_{i=1}^{n}\{|\hat{p}_{i}-p_{i}|\geq\delta p_{i}\}\right)\leq 2n\exp\left(-\frac{\delta^{2}e^{-S_{\rm max}}k}{3}\right),$ (23)
so that finally,
$2n\exp\left(-\frac{\delta^{2}e^{-S_{\rm max}}k}{3}\right)<\alpha\quad\Longrightarrow\quad k>\frac{3e^{S_{\max}}}{\delta^{2}}\log\left(\frac{2n}{\alpha}\right)$
Appendix C Proofs
C.1 Proof of Lemma 3.4
Since we assume that $S(X)$ is a continuous variable with density with respect to the Lebesgue measure, it follows from the properties of IT scores that it is distributed according to the density $p(S(X)=s)=e^{-s}$ for $s\geq 0$. Accordingly, the conditional distribution of $S(X)$, given $S(X)\in[s_{X},S(y)]$ has the density $p(S(X)=s\mid S(X)\in[s_{X},S(y)])=e^{-s}/(e^{s_{X}}-e^{-S(y)})$. We want to compare averages of the expressions
$P(S(Y)\geq S(y)|S(X)=s)\quad\hbox{ versus }\quad e^{-|S(y)-s|_{+}}$
over the interval $s_{X},S(y)]$. Averaging the difference of the two expressions with respect to the above-mentioned density yields
$\int_{s_{X}}^{S(y)}\left(P(S(Y)\geq S(y)|S(X)=s)-e^{-|S(y)-s|_{+}}\right)p(s\mid S(y)\geq S(X)\geq s_{X})ds$
$=P(S(Y)\geq S(y)|S(y)\geq S(X)\geq s_{X})-\int_{s_{X}}^{S(y)}e^{s-S(y)}\frac{e^{-s}}{e^{-s_{X}}-e^{-S(y)}}ds$
$\leq\frac{e^{-S(y)}}{e^{-s_{X}}-e^{-S(y)}}-\int_{\tilde{s}_{X}}^{S(y)}\frac{e^{-S(y)}}{e^{-\tilde{s}_{X}}-e^{-S(y)}}ds=\frac{e^{-S(y)}}{e^{-s_{X}}-e^{-S(x)}}\left(1-S(y)+s_{X}\right)\leq 0.$
The first inequality holds because $P(B|A)\leq P(B)/P(A)$ for any two events $A,B$ and the second because we have assumed $s_{X}\leq S(y)-1$.
Applying Markov’s inequality, we therefore have
$P_{X\mid S(X)\in[s_{X},S(y)]}[P(S(Y)\geq S(y)\mid S(X))\geq c\cdot e^{-|S(y)-S(X)|_{+}}]\leq 1/c.$ (24)
As $X$ may in general be multivariate, $S$ will not in general be injective, and so we cannot substitute conditioning on a particular value of the score for conditioning on the associated $x$, which is required to give a bound on the conditional score. As such, note that
$P(S(Y)\geq S(y)\mid S(X)=s)=\int P(S(Y)\geq S(y)\mid X=x)\,\,p(x\mid S(X)=s)\,dx,$
so that applying Markov’s inequality for any given $s$,
$P_{X\mid S(X)=s}\left[P(S(Y)\geq S(y)\mid X)\geq c\cdot P(S(Y)\geq S(y)\mid S(X))\right]\leq 1/c.$ (25)
As the inequality holds for all $s$, we can take the average of the LHS with respect to
$p(S(X)=s\mid S(X)\in[s_{X},S(y)])$ to say,
$P_{X\mid S(X)\in[s_{X},S(y)]}\left[P(S(Y)\geq S(y)\mid X)\geq c\cdot P(S(Y)\geq S(y)\mid S(X))\right]\leq 1/c.$ (26)
Applying the union bound to the events in equations 24 and 26, the joint event
$P(S(Y)\geq S(y)\mid S(X))\leq c\cdot e^{-|S(y)-S(X)|_{+}}\quad{\rm and}$
$P(S(Y)\geq S(y)\mid X)\leq c\cdot P(S(Y)\geq S(y)\mid S(X))$
occurs with probability at least 12/c1-2/c, and in such case the following inequality holds:
$P(S(Y)\geq S(y)\mid X=x)\leq c^{2}\cdot e^{-|S(y)-S(x)|_{+}}.$ (27)
Finally, noting we can exchange $S(Y)\geq S(y)$ for $\tau(Y)\geq\tau(y)$ due to the properties of IT scores, and taking the negative logarithm of both sides, for $x$ sampled at random according to $P(X=x\mid S(X)\in[s_{X},S(y)])$,
$S(y\mid x)\geq|S(y)-S(x)|_{+}-2\log c,$ (28)
holds with probability at least 12/c1-2/c, as desired.
C.2 Proof of Lemma 3.5
$P(\tau(Y)\geq\tau(y)\mid X=x)=P(S(Y)\geq S(y)\mid X=x)=P(S(Y)\geq S(y)\mid S(X)=S(x))$
$\leq P(S(Y)\geq S(y)\mid S(X)\geq S(x))\leq\frac{P(S(Y)\geq S(y))}{P(S(X)\geq S(x))}=\frac{e^{-S(y)}}{e^{-S(x)}}.$
Where the first equality follows from the definition of IT scores and the second from injectivity. The first inequality follows from monotonicity, and the second because $P(B\mid A)\leq P(B)/P(A)$ for any two events $A,B$. The final equality again follows from the definition of IT scores. Taking the negative logarithm of both sides, and remembering that anomaly scores are non-negative, we prove the lemma.
C.3 Proof of Lemma 3.7
Assume without loss of generality that $X_{1},\dots,X_{n}$ is a topological ordering of the DAG. Then $S(X_{i}|X_{1},\dots,X_{i-1})=S(X_{i}|PA_{i})$ holds due to the local Markov condition. Since $S(X_{j}|PA_{j}=pa_{j})$ has density $e^{-s}$ for all $pa_{i}$, also $S(X_{i}|X_{1}=x_{1},\dots,X_{i-1}=x_{i-1})$ has density $e^{-s}$ for all $x_{1},\dots,x_{i-1}$. Hence, $S(X_{i}|PA_{i})$ is independent of $X_{1},\dots,X_{i-1}$.
C.4 Proof of Lemma 3.8
From Lemma 3.7, and the properties of IT scores, if $H_{0}^{j}$ holds then $S(X_{i}\mid PA_{i})$ for $i\neq j$ are independent, standard Exponential random variables. The sum of n1n-1 independent, standard Exponential random variables is Erlang distributed Ibe (2013) with CDF
$F(x)=1-e^{-x}\sum_{l=0}^{n-2}\frac{x^{l}}{l!},$
and as such, $S_{\rm sum}$ is Erlang distributed. By the probability integral transform $U:=F(S_{\rm sum})$ is uniformly distributed on $[0,1]$, and so by symmetry, so is 1U1-U. Again by the probability integral transform, $\log(1-U)$ is standard Exponential, so finally noting that $S=-\log(1-F(S_{\rm sum}))$, we have the result.
C.5 Proof of Theorem 3.9
The theorem is a direct result of Lemma 3.7 and Lemma 3.8.
C.6 Proof of Theorem 3.11
$s_{\rm sum}:=\sum_{i\neq j}S(x_{i}|pa_{i})\geq\sum_{i\neq j}|S(x_{i})-S(pa_{i})|_{+}=:\hat{s}_{\rm sum},$ (29)
where the inequality follows from score typicality of event $(x_{1},\dots,x_{n})$, and $s_{\rm sum}$ and $\hat{s}_{\rm sum}$ are the realisations of $S_{\rm sum}$ and $\hat{S}_{\rm sum}$, respectively. The theorem then follows directly from Theorem 3.9 with $\hat{S}_{\rm sum}$ in place of $S_{\rm sum}$.
C.7 Proof of Theorem 3.12
From Lemma 3.7 we have that the conditional anomaly score for the anomaly with the maximum score gap must be at least $\delta_{\rm max}$. From Lemma 3.4, and the properties of IT anomaly scores, the conditional anomaly scores are independent and each distributed according to density $p(s)=e^{-s}$. Under $H_{0}^{\mathrm{max}}$, assume without loss of generality that $x_{n}$ corresponds to the true root cause. Then the remaining (n1)(n-1) conditional scores are i.i.d. $\mathrm{Exp}(1)$, and
$P\big(S(X_{i}\mid\mathrm{PA}_{i})<\delta\text{ for all }i<n\big)=\big(P(S(X_{i}\mid\mathrm{PA}_{i})<\delta)\big)^{n-1}=(1-e^{-\delta})^{n-1}.$
Therefore, the probability of observing at least one conditional anomaly score of at least $\delta_{\max}$ among the (n1)(n-1) non-root causes is $1-(1-e^{-\delta_{\max}})^{n-1}$. This probability upper bounds the probability of observing at least one score gap of $\delta_{\max}$ or greater as score gaps lower bound conditional scores.
C.8 Proof of Theorem 3.13
The proof of the theorem is already almost complete, following the argument in the main text. As noted there, $H_{0}^{\text{top-}k}$ implies that along the causal path the score of $x_{\pi(1)}$ must be higher than an ancestral node by at least $\Delta_{k}$. Lemma 3.3 along with the coarsening argument, implies that, individually, the hypothesis that the mechanism generating $x$ worked as expected, where $S(x)$ is higher than the score of an ancestral node by $\Delta_{k}$, can be rejected at level $p\leq e^{-\Delta_{k}}$. Therefore, along the causal path, the probability that the score of any of the (maximum of) $n$ nodes exceeds an ancestral node by $\Delta_{k}$ is at most $n\cdot e^{-\Delta_{k}}$, applying the union bound. However, note that we need to reapply this argument for each of the chains incident on $X_{\pi(1)}$ in the graph. The number of incident chains is at most $d_{\rm max}$, and so again applying the union bound, we arrive at the final bound given in the theorem.
C.9 Proof of Proposition A.1
With the event $E$ as in Eq. 3 and $C^{\pi}(j):=C(j|I^{\pi<j})$ we have
$C^{\pi}(j)$ $=\log\frac{P(E|{\bf N}_{\pi<j~\cup\left\{j\right\}}={\bf n}_{\pi<j~\cup\left\{j\right\}})}{P(E|{\bf N}_{\pi<j}={\bf n}_{\pi<j})}\overset{i}{=}\log\frac{P(E|{\bf X}_{\pi<j~\cup\left\{j\right\}}={\bf x}_{\pi<j~\cup\left\{j\right\}})}{P(E|{\bf X}_{\pi<j}={\bf x}_{\pi<j})}$
$\overset{ii}{=}\log\frac{P(E|do({\bf X}_{\pi<j~\cup\left\{j\right\}}={\bf x}_{\pi<j~\cup\left\{j\right\}})}{P(E|do({\bf X}_{\pi<j}={\bf x}_{\pi<j}))}.$ (30)
(i) and (ii) are seen as follows:
$P(E|{\bf N}_{\mathcal{I}})=P(E|{\bf X}_{\mathcal{I}}))=P(E|do({\bf X}_{\mathcal{I}})).$ (31)
The first equality in Eq. 31 follows from $X_{n}\perp X_{\mathcal{I}}\,|{\bf N}_{\mathcal{I}}$ and because $\bf X$$_{\mathcal{I}}$ is a function of $\bf N$$_{\mathcal{I}}$. The second one follows because conditioning on all ancestors blocks all backdoor paths. Note that since $\pi$ is a topological ordering of the nodes, all $\pi<j$ are ancestors of $j$.
C.10 Proof of Lemma A.2
Denote $q({\mathcal{S}})=P(E\,|\,{\bf N}_{{\mathcal{S}}}={\bf n}_{{\mathcal{S}}})$ then we have that
$C(\mathcal{R}|{\mathcal{S}})$ $=\log q({\mathcal{S}}\cup\mathcal{R})-\log q({\mathcal{S}})$
$=\left(\log q({\mathcal{S}}\cup\mathcal{R})-\log q({\mathcal{S}}\cup\mathcal{R}\setminus\left\{j_{k}\right\})\right)+\ldots$
$\left(\log q({\mathcal{S}}\cup\mathcal{R}\setminus\left\{j_{k}\right\})-\log q({\mathcal{S}}\cup\mathcal{R}\setminus\left\{j_{k},j_{k-1}\right\}\right))+\ldots$
$+\left(\log q({\mathcal{S}}\cup\left\{j_{1}\right\})-\log q({\mathcal{S}})\right)=C(j_{1}|{\mathcal{S}})+\sum_{i=2}^{k}C(j_{i}|{\mathcal{S}}\cup\{j_{1},\dots,j_{i-1}\}).$
C.11 Proof of Proposition A.3
By definition, we have that
$C(\mathcal{R}|{\mathcal{S}})=\log\frac{P(E|{\bf N}_{R}={\bf n}_{R},{\bf N}_{S}={\bf n}_{S})}{P(E|{\bf N}_{S}={\bf n}_{S})},$
and we use $P(E|{\bf n}_{S})$ instead of $P(E|{\bf N}_{S}={\bf n}_{S})$ when it is clear from the context.
Further, $C(\mathcal{R}|{\mathcal{S}})$ is actually a function of ${\bf n}_{R}$ and ${\bf n}_{S}$. For fixed ${\bf n}_{S}$, define the set $B:=\{{\bf n}_{R}\,|C(\mathcal{R}|{\mathcal{S}})\geq\alpha\}$. It can equivalently be described by
$B=\{{\bf n}_{R}\,|\log\frac{P(E|{\bf n}_{R},{\bf n}_{S})}{P(E|{\bf n}_{S})}\geq\alpha\}=\{{\bf n}_{R}\,|P(E|{\bf n}_{R},{\bf n}_{S})\geq P(E|{\bf n}_{S})\cdot e^{\alpha}\}.$
We thus have
$P(E|{\bf n}_{R}\in B,{\bf n}_{S})\geq P(E|{\bf n}_{S})\cdot e^{\alpha}.$
Hence,
$\frac{P(E,{\bf n}_{R}\in B|{\bf n}_{S})}{P({\bf n}_{R}\in B|{\bf n}_{S})}\geq P(E|{\bf n}_{S})\cdot e^{\alpha}.$
Using
$P(E|{\bf n}_{S})\geq P(E,{\bf n}_{R}\in B|{\bf n}_{S})$
we obtain
$P({\bf n}_{R}\in B|{\bf n}_{S})\leq e^{-\alpha}.$
Appendix D Visualising polytrees
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$
(a) General DAG
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$
(b) Polytree
Figure 2: (a) A general DAG may contain undirected cycles, as between $X_{1}-X_{2}-X_{3}-X_{1}$. (b) A polytree cannot contain undirected cycles so could not have an edge between $X_{1}$ and $X_{2}$ without first removing another.
Appendix E Is increase of scores along causal pathways rare?
Li et al. Li et al. (2024) describe a condition under which the $z^{2}$-score of the effect $X_{j}$ can be larger than the $z^{2}$-score from the cause $X_{i}$ in a DAG with linear structural equations. Their Theorem 2.1 states that this happens (in the limit of large perturbations, see the details further below) if and only if the variances $\sigma_{i}^{2},\sigma_{j}^{2}$ of $X_{i},X_{j}$ satisfy
$\sigma_{j}^{2}<\alpha_{i\to j}^{2}\sigma_{i}^{2},$ (32)
where $\alpha_{i\to j}^{2}$ denotes the structural coefficient for the effect from $X_{i}$ to $X_{j}$ (which consists of the sum over the product of all direct influences along the paths from $i$ to $j$). Although our results refer to IT scores instead of $z^{2}$-scores, it is at the same time an example for increasing IT scores since the $z^{2}$-score can be turned into an IT score by a monotonic recalibration. It is therefore important that we understand this result if we are to understand if we expect SCORE ORDERING to be applicable in cases beyond polytrees. For an intuitive understanding we should first note that (32) describes the regime of strong negative confounding. This is because an unconfounded causal influence $X_{j}=\alpha_{i\to j}X_{i}+E_{j}$ with independent error term $E_{j}$ results in $\sigma_{j}^{2}=\alpha_{i\to j}^{2}\sigma_{i}^{2}+{\rm Var}(E_{j})$. We are therefore in the regime where the noise is so strongly negatively correlated with the cause that it decreases the variance instead of increasing it.
Let us now discuss whether this phenomenon is rare or not when $X_{i}$ and $X_{j}$ are variables in a larger network (note that the following working is strongly inspired by Li et al. (2024)). To this end let $X_{1},\dots,X_{n}$ be causally ordered with structural equations
${\bf X}=A{\bf X}+{\bf N},$ (33)
where $A$ is a strictly lower triangular matrix and $\bf N$} is the vector of independent noise variables. To simplify notation, let all variables have zero mean.
The covariance matrix $\Sigma_{{\bf XX}}$ can be derived by well-known algebra:
$\Sigma_{{\bf XX}}=(I-A)^{-1}\Sigma_{{\bf NN}}(I-A)^{-T},$
where $\Sigma_{{\bf NN}}$ denotes the covariance matrix of the noise, which is diagonal by assumption. Defining
$L:=(I-A)^{-1}\Sigma_{{\bf NN}}^{1/2},$
we obtain the unique Cholesky decomposition of $\Sigma_{{\bf XX}}$ since $\Sigma_{{\bf XX}}=LL^{T}$, where $L$ is lower diagonal (with non-zero diagonal by construction if we assume non-zero noise variance). The matrix $L$ has interesting interpretations:
(i) $L$ generates the observations from independent noise variables as sources, formally ${\bf X}=L\tilde{{\bf N}}$, where $\tilde{{\bf N}}:=\Sigma_{{\bf NN}}^{-1/2}{\bf N}$ is the standardized version of the noise from the original structural equation (33).
(ii) The squared row sums of $L$ thus show the variance of the observed variables, that is, $\sum_{j}L^{2}_{ij}=\sigma_{i}^{2}$. Henceforth, we will assume it to be $1$ without loss of generality.
(iii) The $i$th column of $L$ describes how shifting the noise $\tilde{n}_{i}$ to $\tilde{n}_{i}+1$ changes the observations $\bf x$}, that is, the observation ${\bf x}=L\tilde{{\bf n}}$ is changed to ${\bf x}+Le_{i}$, where $e_{i}$ denotes the $i$th canonical basis vector.
Since all the $X_{j}$ are standardized and centered, their $z^{2}$ score is simply their squared value $x_{j}^{2}$. If we shift the noise $\tilde{n}_{i}$ by a large value, the $z^{2}$-scores $x^{2}_{i}$ of the root cause $X_{i}$ versus the score $x_{j}^{2}$ of the effect $X_{j}$ satisfies the ratio $L_{ii}^{2}/L_{ij}^{2}$ in the limit of infinitely strong perturbation. The question of whether scores of effects are typically smaller than scores of their root causes now boils down to the following question: Given a lower triangular matrix $L$ whose row vectors have norm $1$. Is $L^{2}_{ij}$ “typically” smaller than $L^{2}_{ii}$? We will discuss this question for the assumption that $n$ is large and that $\rho:=\min_{i}\{L^{2}_{ii}\}$ is much larger than $1/n$. This assumption simply formalizes the idea that no $X_{i}$ is close to be a deterministic function of its ancestors. It may seem strong, but we could easily extend the below discussion to the case where only most of the variables satisfy this condition. [^8] Note that simulated data can easily violate this assumption since many simulation schemes result in artifacts where variables that are late in the causal order can be almost perfectly reconstructed from others, a phenomenon called $R^{2}$-sortability in Reisach et al. (2023). We conclude:
$\sum_{1\leq i<j}L^{2}_{ij}\leq 1-\rho.$
Hence, the number $k$ of $i$ with $i<j$ for which $L^{2}_{ii}<L^{2}_{ij}$ is at most $(1-\rho)/\rho$. Using $\rho\gg 1/n$ we conclude $k\ll n$. Thus, only for a small fraction $k/n$ of root causes $i$, the score $L^{2}_{ii}$ of the root cause is smaller than the score $L^{2}_{ij}$ of the downstream effect $j$. In this sense, we may consider effects with larger score than the root cause as a rare phenomenon, even for fixed causal structural equations, not only as a statistical statement over multiple randomly generated structures.
Appendix F Experimental details and further experiments
F.1 Experimental details
To run Circa and Counterfactual, we used the implementations available in Hardt et al. (2024) using default parameters. We wrote our own implementation of Traversal (with anomaly score threshold of 3), and our own algorithms, SMOOTH TRAVERSAL AND SCORE ORDERING. The code for Cholesky is available at Li et al. (2024), however, there are in fact three different algorithms. For all experiments involving real-world data (see subsection F.6 below) we applied their "main" algorithm, using default parameters. For the synthetic experiments, we had to opt for the "highdim" version as the alternatives had already become prohibitively slow for the graph sizes considered. To generate an SCM for the experiments in Section. 4 (see Fig. 1), we first uniformly sample between 10 and 20 root nodes (20% to 40% of the total nodes of the graph) and uniformly assign to each either a standard Gaussian, uniform, or mixture of Gaussians as its noise distribution. As a second step, we recursively sample non-root nodes. Non-root nodes need not be sink nodes. The number of parent nodes that each non-root node is conditioned on is randomly chosen following a distribution that assigns a lower probability to a higher number of parents. In total, the causal graph is composed of 50 nodes. The parametric forms of the structural equations are randomly assigned to be either a simple feed-forward neural network with a probability of 0.8 (to account for non-linear models) and a linear model. The feed-forward neural network has three layers (input layer, hidden layer, and output layer) where the hidden layer has a number of nodes chosen randomly between 2 and 100. All the parameters of the neural network are sampled from a uniform distribution between -5 and 5. For the linear model, we sample the coefficients of the linear model from a uniform distribution between -1 and 1 and set the intercept to 0. In both cases, we use additive Gaussian noise as the relation between the noise and the variables.
To generate data for the non-anomalous regime, we sample the noise of each of the variables and propagate the noise forward using the previously sampled structural equations. As mentioned in the main text, to produce anomalous data, we choose a root cause at random from the list of all nodes and a target node from the set of its descendants (including itself). Then we sample the noise of all variables and modify the noise of the root cause by adding $x$-times standard deviations (of its marginal distribution), where $x\in\{2,2.1,2.2,\dots,3\}$, and propagate the modified noise through the SCM to obtain a realisation of each variable. We repeat this process 100 times for each $x$ value added to the standard deviation and consider the algorithm successful if its result coincides with the chosen root cause.
Computing Infrastructure. All the experiments were run on a MacBook Pro with 16GB of memory with an Apple M1 processor.
Figure 3: Runtimes of the algorithms for the experiment in Fig. 1; that is 50 nodes (left), an SCM with 100 nodes (center) and one with 1k nodes (right) (refer to the generation process above). The boxplots are produced using the default implementation in Matplotlib Hunter (2007). Note the log scale in the vertical axis.
F.2 Further experiments varying root cause anomaly strength
Methods which require multiple samples from the anomalous regime
Most RCA approaches require multiple samples from the anomalous regime. For completeness, we extended the evaluation in Figure 1 to include two such methods: RCD Ikram (2023) and $\varepsilon$-Diagnosis Shan et al. (2019), as neither require the causal graph (see Figure 4). For both methods we made use of the implementations available in the PyRCA library Liu et al. (2023) using default parameters. While all other methods were provided with only a single sample from the anomalous regime, both RCD and $\varepsilon$-Diagnosis were provided with 100, and so we believe that direct comparison is not appropriate. The experiment was otherwise identical to that described above in Appendix F. Nonetheless, we see that while RCD has very strong performance, with a top-1 recall close to one at all anomaly strengths, $\varepsilon$-Diagnosis is the opposite, with a recall close to zero for all anomaly strengths. It is worth noting that the original authors of the RCD method Ikram et al. (2022) show in their supplemental material that performance is very poor in the low-sample regime (<100 anomalous samples, see Figure 2 therein). We conclude that RCD could be preferable in scenarios where one has access to many samples in the anomalous period, but that in the low-sample regime SCORE ORDERING is preferable.
Figure 4: True positive rate for identifying the root cause against anomaly strength injected at the root cause, including RCD and $\varepsilon$-Diagnosis. Note that both RCD and $\varepsilon$-Diagnosis are given 100 samples from the anomalous period, while all other methods are given only one.
How do the algorithms perform when the structural causal model is linear?
To test whether the performance of Circa and Cholesky in Fig. 1 was poor compared to the other algorithms due to the assumption of linearity, we reproduced the experiment as detailed in Section 4 and Appendix F above, with the only change being that now all structural equations were sampled to be linear. Indeed, Fig. 5 shows that Circa now becomes one of the best performing algorithms, alongside SMOOTH TRAVERSAL, Traversal, and Counterfactual. Similarly, although Cholesky remains below the top performers, its performance is notably improved at all anomaly strengths, even now outperforming SCORE ORDERING.
Figure 5: True positive rate for identifying the root cause against anomaly strength injected at the root cause, when all structural equations are linear.
How do the algorithms perform when the causal graph is a polytree?
In all synthetic experiments, we placed no restrictions on the structure of the causal graph other than that it be a DAG. Our main theorems are stated, however, for polytrees. We therefore repeated the experiments from Figure 1 with the single change constraining the simulated causal graphs to be polytrees (see Figure 6). Unsurprisingly, methods which do not make use of the causal graph have unchanged performance, and and those that do are either unchanged or show improvement, including SMOOTH TRAVERSAL. Interestingly, Circa shows a marked improvement in performance despite not depending on the polytree assumption.
Figure 6: True positive rate for identifying the root cause against anomaly strength injected at the root cause, when all causal graphs are polytrees. Note that RCD and $\varepsilon$-Diagnosis are given 100 samples from the anomalous period, while all other algorithms are given only one.
F.3 Further experiments by varying the number of nodes in the graph
We also run experiments by fixing the number of standard deviations added ($x$ in Section. 4) to 3 and varying the number of nodes in the SCM. Here, we have chosen the numbers $\{20,40,60,80,100\}$. We observe in Fig. 7 that the performance for Traversal, SMOOTH TRAVERSAL, and Counterfactual does not change much for different graph sizes, whereas the performance of SCORE ORDERING, Cholesky and Circa decreases slightly for larger graph sizes.
Figure 7: True positive rate of the compared algorithms in identifying the root cause with increasing numbers of nodes.
F.4 How does performance change when the given causal graph is misspecified
To mimic the scenario where one has access to an imperfect causal graph, we investigate how robust the performance each of the algorithms (which require a causal graph) is to its misspecification (8). The data were generated as described at the beginning of Appendix F, but with a fixed anomaly strength of $x=3.0$, all structural equations linear, and the graph provided to each algorithm randomly altered to introduce mismatches with the true causal graph.
Starting from the sampled, true causal graph, random edge additions, removals and reversals were introduced according to a desired target Structural Hamming Distance Acid and de Campos (2003); Tsamardinos et al. (2006) from the true graph. To control for the density of the graph, an equal number of edges are randomly added as removed. In particular, for true causal graph $\mathcal{G}=(V,E)$, the total number of edges added and removed is sampled uniformly at random from $n_{\rm add/rem}=[\text{SHD}(\mathcal{G},\mathcal{G}_{alt})-|E|,\,\,\text{SHD}(\mathcal{G},\mathcal{G}_{alt})]$, with the number of edges flipped $n_{\rm flip}=\text{SHD}-2\cdot n_{\rm add/rem}$. Of the edges in $\mathcal{G}$, $n_{\rm add/rem}$ are randomly selected for removal, and an equal number of ‘missing’ edges are selected for addition. Of the remaining ‘un-removed’ edges in $\mathcal{G}$, $n_{\rm flip}$ are reversed. If the resulting graph is not a DAG, the whole process is repeated until it is. The procedure for generating a randomly misspecified graph given a ground truth matches the "edge domain expert" procedure described in Eulig et al. (2025).
We chose to sample all linear structural equations as the relationship between top-1 recall and SHD was the same as when sampling non-linear structural equations for SMOOTH TRAVERSAL, Traversal and Counterfactual, but was more noteworthy for Circa. The reason being that the performance for Circa is already so low for non-linear structural equations, that a considerable drop in performance was not easily perceptible.
We observe that all four evaluated algorithms show a considerable drop in performance as the SHD of the given graph from the true causal graph increased, with the most pronounced drop occurring for Circa. However, even when half of all edges are removed/flipped or added elsewhere in the graph (SHD/|E|=1/|E|=1), SMOOTH TRAVERSAL, Traversal and Counterfactual still achieve a top-1 recall of approximately 0.6 to 0.7, showing that all three algorithms are relatively robust to graph misspecification.
Figure 8: Top-1 recall for identifying the root cause using a graph with an increasing SHD (average per edge) from the true causal graph.
F.5 How do the algorithms compare in terms of their practicality?
It is important to keep in mind when interpreting the performances of each of the algorithms their different input requirements. Counterfactual either needs to be provided the SCM or must attempt to learn it from data. Traversal, Circa and SMOOTH TRAVERSAL require the causal graph as input but with differing assumptions on its structure. Traversal makes no assumptions, Circa assumes the SCM is linear and attempts to learn it from data, given the graph, and SMOOTH TRAVERSAL has guarantees only for the case the graph is a polytree (though it may be applied even if this assumption is not met). Traversal, however, requires an additional tuning parameter: the anomaly score threshold, while SMOOTH TRAVERSAL does not have any such free parameter. Both SCORE ORDERING and Cholesky do not even require the causal graph, however, Cholesky assumes the underlying SCM is linear. In synthetic data experiments, SCORE ORDERING consistently outperformed Cholesky other than when the SCM was restricted to be linear (see Fig. 5 above). Cholesky also scales more poorly in terms of run-time (see Fig. 3) due to the need to estimate the covariance matrix. In comparison, SCORE ORDERING makes very weak demands on the input, is efficient, and consistently performs well across experiments.
F.6 PetShop dataset
Hardt et al. Hardt et al. (2024) introduce a dataset from cloud computing, specifically designed for evaluating root cause analyses in microservice-based applications. The dataset is collected from a microservice application and includes 68 injected performance issues, which increase latency and reduce availability throughout the system. Latency and availability are so-called ‘metrics’ of the microservice application, and measure the length of time it takes for a request to the service to be processed, and the fraction of the time a request returns an error, respectively. The dataset also encompasses three traffic scenarios: low, high, and temporal. These correspond to the amount and periodicity of user traffic to the application. This presents a challenging task due to high missingness, low sample sizes, and near-constant variables. Furthermore, the ground truth causal graph is only partially known, as we must treat the edge-inverted call graph as a proxy for the causal graph.
In addition to the approaches evaluated by Hardt et al. (see Hardt et al. (2024) for more details), which we have reproduced below, we evaluated our algorithms in both top-1 recall (Table. 1) and top-3 recall (Table. 2). In addition, we included, as an additional baseline, the Traversal algorithm as described in Section 4. This is added as although the evaluation already includes a ‘Traversal’ algorithm implemented by Hardt et al., here called ‘petshop Traversal’, their algorithm does not match the standard approach as described in Liu et al. (2021); Chen et al. (2014); Lin et al. (2018). Instead, ‘petshop Traversal’ has been designed for use on the PetShop dataset, takes into account particular aspects of that data, and contrary to standard traversal algorithms, scores potential root causes according to anomalousness of whole paths of nodes from root cause to target node. ‘Petshop Traversal’ should not, therefore, be treated as an ‘out-of-the-box’ baseline.
We observe that SMOOTH TRAVERSAL and SCORE ORDERING perform very well in top-1 and top-3 recall, often being the best performing algorithms, and clearly outperforming simple Traversal and Cholesky. However, while SMOOTH TRAVERSAL always matches or exceeds the performance of simple Traversal in top-1 recall, on latency issues it nonetheless tends to perform relatively poorly compared to SCORE ORDERING.
Table 1: Top-1 recall, with ties. Multiple anomalies can receive the same, highest, anomaly score. The root cause being in the top-1 means that it is among the variables receiving the highest score.
graph given graph not given this paper
traffic metric petshop Circa counter- $\epsilon$-diagnosis rcd correlation Traversal Cholesky SCORE ORDERING SMOOTH TRAVERSAL
scenario Traversal factual
low latency 0.57 0.36 0.36 0.00 0.07 0.43 0.14 0.29 1.00 0.14
low availability 0.50 0.42 0.00 0.00 0.58 0.75 0.42 0.00 0.75 0.75
high latency 0.57 0.50 0.57 0.00 0.00 0.64 0.14 0.14 1.00 0.14
high availability 0.33 0.00 0.00 0.00 0.00 0.83 0.33 0.00 1.00 1.00
temporal latency 1.00 0.75 0.38 0.12 0.25 0.62 0.25 0.50 0.75 0.25
temporal availability 0.38 0.38 0.00 0.00 0.50 0.62 0.25 0.00 1.00 1.00
Table 2: Top-3 recall, with ties. Multiple anomalies can receive the same, highest, anomaly score. The root cause being in the top-3 means that it is among the variables ranked in the top three including all variables tied at the top.
graph given graph not given this paper
traffic metric petshop Circa counter- $\epsilon$-diagnosis rcd correlation Traversal Cholesky SCORE ORDERING SMOOTH TRAVERSAL
scenario Traversal factual
low latency 0.57 0.86 0.71 0.00 0.21 0.57 0.14 0.57 1.00 0.86
low availability 1.00 1.00 0.42 0.00 0.75 0.92 0.50 0.00 1.00 1.00
high latency 0.79 1.00 0.86 0.00 0.07 0.79 0.14 0.57 1.00 0.93
high availability 1.00 0.00 0.00 0.33 0.00 0.92 0.33 0.00 1.00 1.00
temporal latency 1.00 1.00 0.50 0.12 0.75 0.75 0.25 0.63 1.00 0.75
temporal availability 1.00 1.00 0.25 1.00 0.12 0.75 0.25 0.00 1.00 1.00
Note, however, the caveat that owing to the small sample size in the normal period in the PetShop dataset, it is frequently the case that multiple anomalies receive the same highest IT anomaly score (see Table 3 for numbers and proportions of nodes with tied scores in eah traffic scenario). In the most severe cases there can be as many as 20 nodes tied with the maximum score (in the high traffic scenario), but it is typically fewer than 10. The multiplicity of highest scores is due to the fact that for the issue at hand all these metrics were more extreme than ever sampled in the normal period. In this case, the root cause being in the top-1 means that it is among the variables receiving the highest score. This problem should become less severe for datasets where anomaly scorers are trained on a longer history. Since root cause analysis without any graphical information is a notoriously hard problem, returning short lists of potential root causes, even if they contain 10 or more candidates, can still result in a drastic reduction of the search space, and so is of vital use in real applications.
Table 3: Average number and proportion of nodes in each traffic scenario and metric with a tied marginal anomaly score
Traffic scenario Metric Average number Average proportion
high traffic Latency 18.26 41.50%
high traffic Availability 16.46 37.41%
low traffic Latency 9.96 23.72%
low traffic Availability 6.54 15.57%
temporal traffic1 Latency 5.75 14.02%
temporal traffic1 Availability 7.00 17.07%
temporal traffic2 Latency 5.75 14.02%
temporal traffic2 Availability 7.12 17.38%
We, nonetheless, also report the top-1 (Table. 4) and top-3 (Table. 5) recall after randomly selecting among the nodes tied with the highest anomaly score. As expected, we observe an approx. 10x decrease in top-1 recall for Traversal, SCORE ORDERING and SMOOTH TRAVERSAL, corresponding to selecting at random from among the approx. 10 candidate root causes with tied highest scores. Similar decreases are present in top-3 recalls for Traversal and SCORE ORDERING whenever the true root cause is among those with the tied highest score. Note, however, that SMOOTH TRAVERSAL generally still has high top-3 recall even after random selection of the top-1 node, outperforming both Traversal, SCORE ORDERING and Cholesky in all but one evaluation. Note that the performance of Cholesky does not considerably drop after random selection of ties, as variables rarely receive the same score.
Table 4: Top-1 recall, with random selection among ties.
this paper
traffic metric Traversal Cholesky SCORE ORDERING SMOOTH TRAVERSAL
scenario
low latency 0.02 0.29 0.10 0.02
low availability 0.06 0.00 0.10 0.10
high latency 0.01 0.14 0.06 0.01
high availability 0.03 0.00 0.06 0.09
temporal latency 0.04 0.50 0.10 0.04
temporal availability 0.04 0.00 0.13 0.14
Table 5: Top-3 recall, with random selection among ties
this paper
traffic metric Traversal Cholesky SCORE ORDERING SMOOTH TRAVERSAL
scenario
low latency 0.02 0.57 0.10 0.73
low availability 0.10 0.00 0.31 0.31
high latency 0.01 0.50 0.06 0.80
high availability 0.03 0.00 0.06 0.09
temporal latency 0.04 0.63 0.18 0.54
temporal availability 0.04 0.00 0.13 0.14
F.7 Sock-shop dataset
Similar to the PetShop dataset described above, Pham et al. Pham et al. (2024) introduces another dataset from cloud computing for evaluating root cause analyses. The dataset includes 125 injected performance issues encompassing five types of common faults: CPU hog - ‘cpu’, memory leak - ‘mem’, disk IO stress - ‘disk’, network delay - ‘delay’, and packet loss - ‘loss’. We again used the edge-inverted call graph for the Sock-shop application as a proxy for the causal graph. We evaluated the performance of our algorithms in top-1 (Table 6) and top-3 (Table 7) recall. It is important to note that while for most algorithms we provided only a single anomalous sample (the first sample from the anomalous period) for RCD and $\epsilon$-Diagnosis we provided all 361 anomaly samples and so direct comparison is not recommended: both algorithms were included rather for completeness of evaluation.
We observe that on three out of five issue types (delay, disk and loss), SCORE ORDERING is the top performing approach for top-1 and top-3 recall. In each instance, SMOOTH TRAVERSAL also performed competitively. For issue types cpu and mem, however, both SCORE ORDERING and SMOOTH TRAVERSAL performed poorly, although performance was comparable among all algorithms which were provided with only a single anomalous sample. Given both RCD and $\epsilon$-Diagnosis performed considerably better for both issues, this would suggest that the effect of the fault was not easily discernible from only a single sample in the anomaly period.
As for PetShop, we also report the top-1 (Table 8) and top-3 (Table 9) recall after randomly selecting among nodes with tied ranks. Here we observe that although the performance of SMOOTH TRAVERSAL is competitive for delay, disk and loss issues, both SCORE ORDERING and SMOOTH TRAVERSAL are outperformed by Circa and/or Cholesky. As both algorithms assume linear causal models, and that both algorithms struggled in non-linear settings in our simulation studies (see Figure 1), this may suggest that the causal relationships in Sock-shop are well modelled by linear relationships. As before, for cpu and mem issues, while all methods which use only a single anomaly sample perform very poorly, RCD performs well.
Table 6: Top-1 recall, with ties. Multiple anomalies can receive the same, highest, anomaly score. The root cause being in the top-1 means that it is among the variables receiving the highest score. These methods were provided with all samples from the anomaly period, whereas all else were given only one.
cpu delay disk loss mem
SCORE ORDERING 0.08 0.80 0.84 0.72 0.08
SMOOTH TRAVERSAL 0.00 0.72 0.76 0.60 0.00
Cholesky 0.04 0.56 0.48 0.40 0.12
Circa 0.04 0.52 0.68 0.64 0.16
Counterfactual 0.04 0.28 0.00 0.12 0.04
Traversal 0.04 0.80 0.76 0.72 0.12
rcd 0.88 0.52 0.52 0.44 0.36
epsilon diagnosis 0.24 0.08 0.16 0.20 0.00
Table 7: Top-3 recall, with ties. Multiple anomalies can receive the same, highest, anomaly score. The root cause being in the top-3 means that it is among the variables ranked in the top three, including tied ranks. These methods were provided with all samples from the anomaly period, whereas all else were given only one.
cpu delay disk loss mem
SCORE ORDERING 0.24 0.92 0.92 0.88 0.16
SMOOTH TRAVERSAL 0.12 0.80 0.84 0.72 0.12
Cholesky 0.16 0.88 0.88 0.84 0.24
Circa 0.28 0.72 0.80 0.80 0.28
Counterfactual 0.24 0.76 0.52 0.64 0.40
Traversal 0.04 0.80 0.76 0.72 0.12
rcd 1.00 0.52 0.56 0.44 0.52
epsilon diagnosis 0.56 0.52 0.28 0.48 0.00
Table 8: Top-1 recall, with random selection among ties. These methods were provided with all samples from the anomaly period, whereas all else were given only one.
cpu delay disk loss mem
SCORE ORDERING 0.08 0.24 0.29 0.26 0.04
SMOOTH TRAVERSAL 0.00 0.49 0.50 0.47 0.00
Cholesky 0.04 0.56 0.48 0.40 0.12
Circa 0.04 0.52 0.68 0.64 0.16
Counterfactual 0.04 0.28 0.00 0.12 0.00
Traversal 0.02 0.41 0.37 0.47 0.02
rcd 0.88 0.56 0.48 0.40 0.36
epsilon diagnosis 0.24 0.08 0.16 0.20 0.00
Table 9: Top-3 recall, with random selection among ties. These methods were provided with all samples from the anomaly period, whereas all else were given only one.
cpu delay disk loss mem
SCORE ORDERING 0.24 0.67 0.70 0.68 0.14
SMOOTH TRAVERSAL 0.12 0.76 0.83 0.72 0.04
Cholesky 0.16 0.88 0.92 0.84 0.24
Circa 0.28 0.72 0.80 0.80 0.28
Counterfactual 0.24 0.80 0.48 0.64 0.28
Traversal 0.04 0.79 0.75 0.72 0.07
rcd 1.00 0.56 0.56 0.44 0.48
epsilon diagnosis 0.56 0.52 0.28 0.48 0.00
F.8 ProRCA
To provide evaluations of the algorithms in a semi-synthetic setting, we used the ProRCA package, introduced in Dawoud and Talupula (2025). ProRCA consists of a hand-crafted model of a retail service, with a known causal graph, allowing synthesis of data from real-world business scenarios, into which different types of anomalies can be introduced.
There are five types of anomalies that can be injected in the process, affecting different variables. We synthesised data following the introduction of four of these anomaly types, as listed in the first column in Table 10, performing 100 replicates of each. We excluded the ‘COGs’ anomaly type, which introduces an anomaly at the variable ‘UNIT_COST’, as for a range of strengths of the anomaly, we were unable to produce changes in the data that were statistically distinct from those in the non-anomalous regime.
We removed from consideration variables in the graph that had plain text values, such as ‘PROMO_CODE’ as these are unsuitable for RCA, but doing so introduced no unmeasured confounding and so does not affect the analysis. For each anomaly, we considered ‘PROFIT’ to be the target variable. ‘PROFIT’ is not a leaf node in the causal graph and the length of the paths between the root cause and the target variable varies between two and three. We applied Circa, Counterfactual, Traversal, and SMOOTH TRAVERSAL algorithms for which we are able to provide (the marginalised) causal graph, and the Cholesky algorithm and SCORE ORDERING, which only requires data from the observational and the anomalous regime. Only one anomalous sample was provided in each case. The resulting Top-1 recall of the algorithms is reported in Table 10.
Table 10: Top-1 recall, with random selection among ties.
graph given graph not given this paper
Type of anomaly Circa Counterfactual Traversal Cholesky SCORE ORDERING SMOOTH TRAVERSAL
ExcessiveDiscount 0.04 0.89 0.30 0.00 0.04 0.04
FulfillmentSpike 0.96 0.98 0.94 0.74 0.94 0.28
ReturnSurge 1.00 0.88 0.97 0.66 0.00 0.18
ShippingDisruption 0.94 0.94 0.94 0.47 0.95 0.27
The results are mixed. SCORE ORDERING performs well for two types of anomaly. In both cases, the methods that require the graph also perform well. However, requiring the graph for the RCA algorithm might be constraining in most real-world scenarios. There are two types of anomaly where SCORE ORDERING does not perform well: In ‘ExcessiveDiscount’ only the Counterfactual method provides a Top-1 recall above 30%, even for those methods requiring the graph as input. In ‘ReturnSurge’, the Cholesky method performs well (66%) and those methods for which the graph is given provide a recall above 85%. In all types of anomalies, SMOOTH TRAVERSAL has a performance below 30%.